OpenStudy (anonymous):

can someone help me simplify 12/6+3 9/16

OpenStudy (anonymous):

simplify... 12/6+3 = 12/9 = 4/3 aka 1and 1/3 9/16 is as simplified as it gets... :D

OpenStudy (anonymous):

are you multiplying them or adding them or something? i am unsure as to the actual problem

OpenStudy (anonymous):

parenthesis would be a very good tool right now :D

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

it could be 12/6 + 3&9/16 which would be 5 and 9/16